OpenStudy (anonymous):

how do you do (3)/(2x)-(1)/(2(x+6))=1 ?

OpenStudy (p0sitr0n):

you need to have a common number in order to add both fractions. when done the steps, it gives \[(x+9)/x(x+6)=1\] you cross multiply, and move all to one side (left) after, you solve the quadratic equation using formula

OpenStudy (p0sitr0n):

so, from here we get \[x+9 = x(x+6)\] \[x+9 - x^2-6x =0\] use this formula to find roots \[x1,x2 = (-b \pm \sqrt{b^2-4ac})/2a\] where a is the term before x^2 where b is term (coefficient) of x and c is without x so we get a = -1 b=-5 c=9 if theres 2 roots = 2 points of intersection of the parabola with the line y=1 1 root = only 1 negative number under root = no intersection , no solution

OpenStudy (p0sitr0n):

-x^2 -5x +9 =0